OpenStudy (anonymous):

integral of sin(2x-4)dx

OpenStudy (anonymous):

The antiderivative of sin(2x-4) is -1/2 cos(2x-4)

OpenStudy (anonymous):

\[\int\limits_{}^{}dx \sin(2x-4)=\int\limits_{}^{}(1/2) du \sin(u)=-(1/2) \cos(u) = -(1/2) \cos(2x-4)\]

OpenStudy (anonymous):

-1/2cos(2x-4)